The profit that the vender makes per day by selling x pretzels is given by the function g(x)=-.004x^2+302x-500. Find the number of pretzels that must be sold to maximize profit.
Answer by stanbon(48530)

The profit that the vender makes per day by selling x pretzels is given by the function g(x)=-.004x^2+302x-500.
Find the number of pretzels that must be sold to maximize profit.
----------------
Max occurs when x = -b/2a = -302/(2*-0.004)) = 37,750
===========================
Cheers,
Stan H.